OpenStudy (anonymous):

Determine the sum: 25/72 + 25/90 + 25/110 + .......... + 25/9900

OpenStudy (vijay):

112 + 100 = 212

OpenStudy (anonymous):

am i reading it wrong? im seeing: \[\frac{25}{72}+\frac{25}{90}+\frac{25}{110}+\cdots +\frac{25}{9900}\] since each number is less than one, i dont think the answer can be as big as 212.

OpenStudy (anonymous):

do you see a pattern in the denominators Estudier? i got nothing so far >.<

OpenStudy (anonymous):

Cant even tell how many terms there are.

OpenStudy (anonymous):

I suppose 18,20..but this just speculation.

OpenStudy (anonymous):

yeah, i dont see anything other than that >.<

OpenStudy (anonymous):

Lot's of 9's about but the 110 ..

OpenStudy (anonymous):

They all reduce by 5's except the first...

OpenStudy (anonymous):

Maybe it should be 25/70

OpenStudy (anonymous):

that would make sense. These types of problems arent supposed to be this hard >.< lol

OpenStudy (anonymous):

That would do it,I think...

OpenStudy (anonymous):

Blast, u can't get to 9900 with +4k...

OpenStudy (anonymous):

theres gotta be a typo somewhere. It should be clear what type of series we have and how to solve.

OpenStudy (anonymous):

Agree...

OpenStudy (anonymous):

Reckon it's supposed to be 6...to 25/9910 with 25/70 to start. http://www.wolframalpha.com/input/?i=sum+5%2F%2814%2B4k%29%2C+k%3D+1..492

OpenStudy (anonymous):

lolol >.< ridiculous

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Nah, no typo I typed it out myself right off the sheet.

OpenStudy (anonymous):

Best answer your are gonna get...about 6

OpenStudy (anonymous):

Thanks both of you.